# Pseudodifferential damping estimates and stability of relaxation shocks

### v2

@@ 5.1. Kreiss’ glancing
for $\tilde{\gamma}$ bounded and $\varepsilon$ sufficiently small. For the general case accounting for neglected terms, see, e.g., [23, §4], [37], [41, Appendix A], or [18, §5].
@@ 5.2. Variable coefficients and Airy-type turning points
For $a^{\prime}(0)=1$, the point $x=0$ in (5.6), corresponding to the special point $x_{*}$ in the original equations, under the assumptions (5.7) can be recognized as an “Airy-type” turning point; cf. [52, 26]. That is, glancing points in $\lambda$ are replaced for variable coefficients by turning points in $x$. Note in (5.6) that $\Im\lambda$ has been absorbed in shifts of $a$, $x$, hence $\tilde{\tau}$ in (5.2) effectively replaced by $a(x)/\varepsilon$. An important difference from the constant-coefficient case is that $ia(x)/\varepsilon$ is more singular than the former term $\tau$, requiring additional care in the symmetrizer construction. In particular, the classical construction of Kreiss can be expected a priori to succeed only in a small neighborhood of $x=0$, where $|a(x)|/\varepsilon$ is small compared to $1/\varepsilon$. (The setting of Kreiss, translated to our coordinates, is $\tau=O(1)$; what is needed for his arguments is $\tau\ll 1/\varepsilon$ [23, §4].)
@@
(5.10) $R=\begin{pmatrix}1&c\\ 1&-c\end{pmatrix}\qquad L=R^{-1}=(1/2)\begin{pmatrix}1&1/c\\ 1&-1/c\end{pmatrix}$,
@@
(5.16) $S_{+}:=\begin{pmatrix}0&1+i\varepsilon\tilde{\gamma}/2|a|\\ 1-i\tilde{\gamma}/2|a|&0\end{pmatrix}$
@@ 6. A basic large-amplitude damping result in multi-D
$\Re(S_{\alpha}G)=\begin{pmatrix}0&1+\varepsilon\tilde{\gamma}/2\alpha\\ 1-\varepsilon\tilde{\gamma}/2\alpha&0\end{pmatrix}$.
@@
1. The reduction to traceless form (6.6) may be regarded as a variant of [23, Lemma 4.3] in the special case of a Jordan block of dimension $s=2$, designed to highlight the connection with Airy’s equation and the analyses of [8, 26, 54]. As can be seen by the details of the argument, form (5.6) is not needed on the whole line, but only in a vicinity of the turning point $x_{*}$, allowing treatment of more general cases with further care.
2. The crucial construction of the family of symmetrizers $S_{\alpha}$ likewise reflects the key observation of [23, Lemma 4.4] for Jordan blocks of arbitrary dimension $s\geq 2$ that there is a degree of freedom in the choice of symmetrizers allowing arbitrarily large constants in coordinates $2,\dots,s$, allowing absorbtion of remaining error terms.
3. The block structure condition for Airy point at infinity, though phrased here as an additional assumption, is in the hyperbolic constant multiplicity case a consequence of high-frequency dissipativity of the end states, which is implied by our standard assumption of high-frequency resolvent bound (1.3). See [23, Lemma 2.7, Eqn. (2.21)] in the stictly hyperbolic case, [38] in the general constant-multiplicity case.
4. Similarly, augmented consistent splitting at finite Air points may be shown to follow from high-frequency resolvent assumption (1.3), via the high-frequency spectral stability it implies. We believe, but have not shown, that, similarly, the block structure assumption for finite Airy points is forced by (1.3). For the analytic-coefficient case, it is a consequence of the spectral analysis of [26], via exact conjugation to the Airy equation. Analytic approximation of coefficients together with continuity of resolvent bounds seems a promising direction for a proof in the general case, bypassing spectral stability to appeal directly to (1.3).
@@ 8.1. Extensions and simplification
Our results yield immediately a number of simplifications and extensions for existing work. For example, In the 1-D large-amplitude relaxation case, they recover and somewhat simplify the damping result of [34], removing contortions associated with translating linear algebraic estimators to Friedrichs symmetrizer form. Likewise, in the viscous shock case [56, 57], global multi-D stability is known only for hyperbolic-parabolic systems of “Kawashima type” (roughly, simultaneously symmetrizable) at the endstates of the shock. Yet, high-frequency bounds (1.3) have been shown in the different context of bounded-time vanishing viscosity already in [15, 16] by Kreiss symmetrizer techniques under the much weaker, apparently sharp assumption [15, 16, (H5)] that the endstates be stable as constant solutions, a condition on the dispersion relation of the associated Fourier symbol. As simultaneous symmetrizability was used in [56, 57] only to establish a nonlinear damping estimate, we thus extend immediately the multi-D long-time stability results of [56, 57] to the larger class of systems considered in [15, 16], putting the two theories (long-time vs. small-viscosity) finally on par.
@@ 8.3. The large-amplitude multi-D case
As regards our main motivating example, of large-amplitude multi-dimensional shocks in the Saint Venant model, the preliminary WKB analysis has been already done in [54, §2.4, p. 32] for the linear problem, yielding the high-frequency linearized spectral stability needed for our analysis, and verifying the assumptions made in Theorem 6.2. We note further, as pointed out in [54], that Airy points do appear for waves of all amplitudes, both smooth and discontinuous type, so that the analysis of Theorem 6.2 is indeed needed to obtain nonlinear damping. Combining these ingredients with the linearized resolvent bounds obtained in [24] for arbitrary-amplitude smooth multi-D profiles of general relaxation systems, one obtains the result that spectral stability in the sense of a standard Evans function condition [24, 54] implies linearized and nonlinear stability with standard decay rates matching those of the parabolic case [56, 57, 58]. Evans stability has been demonstrated for all such waves by nonrigorous numerics in [54]. It would be very interesting to convert this to rigorous computer-aided proof.
@@ 9.1. Open problems
A further issue is the treatment of discontinuous “subshock”-containing solutions, disregarded in the body of the paper. Such solutions were treated in 1-D in [53], using the 1-D property that all stable boundary conditions are maximally dissipative in a suitably chosen coordinate frame, giving favorable boundary terms in the Friedrichs symmetric energy estimate. In multi-D, of course, this property does not hold, and this is one of the main difficulties in the multi-dimensional theory of hyperbolic boundary value problems (BVP) and shock theory, solved by Kreiss using frequency-dependent estimates effectively reducing to 1-D.
An additional ingredient of the Kreiss theory, as described, e.g., in [40], is the construction of a family of “$K$-symmetrizers,” with arbitrarily large weight $K$ on “stable” (i.e., decaying) modes. These become for any given stable boundary condition eventually maximally dissipative as the parameter $K$ is taken sufficiently large, thus effectively removing the details of the boundary condition. A very interesting open problem is to construct families of $K$-symmetrizers in the present variable-coefficient setting, thereby establishing nonlinear multi-D stability of relaxation profiles with subshocks. We regard this as the next main open problem in the theory.
Note that we did effectively construct $K$-symmetrizers (parametrized by $\alpha\to 0$) in the vicinity of turning points in our treatment of the smooth profiles. In the oscillatory region treated by decomposition into scalar modes, they may be accomplished by directly weighting the desired mode. In the exponential region, they are automatic, as the order $1/\varepsilon$ spectral gives arbitrarily large weight to both modes as $\varepsilon\to 0$. Thus, the problem seems mainly one of matching of/interpolation between these various different regimes.
In short, beyond the various applications described in Section 8, the main contribution of the present work in our view is to take the theory to the threshold of a Kreiss symmetrizer type construction for nonlinear damping of large-amplitude discontinuous profiles in multi-D, under the assumptions of noncharacteristic hyperbolicity (A1) and constant multiplity (A2). Where constant multiplicity fails, e.g., some cases in MHD and anisotropic wave motion, further analysis is needed, even in the constant-coefficient case; see related discussion in [42, 17]. Likewise, failure of noncharacteristicity is an important direction for further study, with applications to periodic roll wave solutions [19, 54]. For, as shown in [49], failure of condition (A1) does not imply failure of damping, nor linear equivalence, but only of the general arguments developed above. Given the important example of Saint Venant roll waves [19, 49], it would be very interesting to investigate this case further using other techniques.

### v3

@@ 5.1. Kreiss’ glancing
for $\tilde{\gamma}$ bounded and $\varepsilon$ sufficiently small. For the general case accounting for neglected terms, see, e.g., [24, §4], [38], [42, Appendix A], or [19, §5].
@@ 5.2. Variable coefficients and Airy-type turning points
For $a^{\prime}(0)=1$, the point $x=0$ in (5.6), corresponding to the special point $x_{*}$ in the original equations, under the assumptions (5.7) can be recognized as an “Airy-type” turning point; cf. [53, 27]. That is, glancing points in $\lambda$ are replaced for variable coefficients by turning points in $x$. Note in (5.6) that $\Im\lambda$ has been absorbed in shifts of $a$, $x$, hence $\tilde{\tau}$ in (5.2) effectively replaced by $a(x)/\varepsilon$. An important difference from the constant-coefficient case is that $ia(x)/\varepsilon$ is more singular than the former term $\tau$, requiring additional care in the symmetrizer construction. In particular, the classical construction of Kreiss can be expected a priori to succeed only in a small neighborhood of $x=0$, where $|a(x)|/\varepsilon$ is small compared to $1/\varepsilon$. (The setting of Kreiss, translated to our coordinates, is $\tau=O(1)$; what is needed for his arguments is $\tau\ll 1/\varepsilon$ [24, §4].)
@@
(5.10) $R=\begin{pmatrix}1&1\\ c&-c\end{pmatrix}\qquad L=R^{-1}=(1/2)\begin{pmatrix}1&1/c\\ 1&-1/c\end{pmatrix}$,
@@
(5.16) $S_{+}:=\begin{pmatrix}0&1+i\varepsilon\tilde{\gamma}/2|a|\\ 1-i\varepsilon\tilde{\gamma}/2|a|&0\end{pmatrix}$
@@ 6. A basic large-amplitude damping result in multi-D
(6.9) $S_{\alpha}=\begin{pmatrix}0&1+i\varepsilon\tilde{\gamma}/2\alpha\\ 1-i\varepsilon\tilde{\gamma}/2\alpha&0\end{pmatrix}$.
@@
1. The reduction to traceless form (6.6) may be regarded as a variant of [24, Lemma 4.3] in the special case of a Jordan block of dimension $s=2$, designed to highlight the connection with Airy’s equation and the analyses of [8, 27, 55]. As can be seen by the details of the argument, form (5.6) is not needed on the whole line, but only in a vicinity of the turning point $x_{*}$, allowing treatment of more general cases with further care.
2. The crucial construction of the family of symmetrizers $S_{\alpha}$ likewise reflects the key observation of [24, Lemma 4.4] for Jordan blocks of arbitrary dimension $s\geq 2$ that there is a degree of freedom in the choice of symmetrizers allowing arbitrarily large constants in coordinates $2,\dots,s$, allowing absorbtion of remaining error terms.
3. The block structure condition for Airy point at infinity, though phrased here as an additional assumption, is in the hyperbolic constant multiplicity case a consequence of high-frequency dissipativity of the end states, which is implied by our standard assumption of high-frequency resolvent bound (1.3). See [24, Lemma 2.7, Eqn. (2.21)] in the stictly hyperbolic case, [39] in the general constant-multiplicity case.
4. Similarly, augmented consistent splitting at finite Airy points may be shown to follow from high-frequency resolvent assumption (1.3), via the high-frequency spectral stability it implies. We believe, but have not shown, that, similarly, the block structure assumption for finite Airy points is forced by (1.3). For the analytic-coefficient case, it is a consequence of the spectral analysis of [27], via exact conjugation to the Airy equation. Analytic approximation of coefficients together with continuity of resolvent bounds seems a promising direction for a proof in the general case, bypassing spectral stability to appeal directly to (1.3).
@@
6.1. Discontinuous profiles: the two cases of Erpenbeck
We now turn to the discontinuous case, assuming the same structural assumptions as in the smooth case– i.e., block-diagonalizability of (6.3) for critical frequencies $\lambda=i/\varepsilon+\gamma$ into scalar and $2\times 2$ blocks, each containing at most one Airy-type turning point, without loss of generality at $x_{1}=0$– but now allowing also a single shock discontinuity at $x_{1}=\tilde{x}$, without loss of generality lying in $\infty,0]$. The case that there is a shock but no turning point is straightforward by the methods of Section 3.4.1. We therefore focus on the case that both shock and turning point occur.
[…]
We discuss each of these cases in turn, describing briefly how the symmetrizer constructions of the previous subsection may be modified to accomodate the shock.
6.1.1. Oscillatory intermediate region
The most delicate case is the oscillatory one, for which the high-frequency Kreiss-Lopatinski encodes information that must be incorporated in the symmetrizer construction. On $(-\infty,\delta)$, $0<\delta\ll 1$, we may use the symmetrizer construction of the smooth case, after the coordinate transformation $x\to-x$. Recall that in diagonalized coordinates, this amounts to the choice $S={\rm diag}\{-1,1\}$ at $x=\delta$. Choosing a symmetrizer on $\delta,\tilde{x}$] matching this value at $x=\delta$ then amounts to solving the bounded interval case treated in Section 4.2, with artifical left boundary condition $u^{+}(0)=u^{-}(0)$.
But, likewise, the WKB study of [27, 55] (encoding Airy dynamics as described in [27, 61], taking the decaying solution at -$\infty$ to the sum of growing and decaying solutions at $0^{+}$) gives a Kreiss-Lopatinsky condition corresponding to the condition (KL) for the bounded interval. Thus, again, the high-frequency resolvent condition implies (KL), and thereby existence of a smooth symmetrizer. To the right of the subshock, there is no turning point, and so uniform exponential dichotomies hold. Thus, we may treat this part by a straighforward weighting as in Section 3.4.1.
6.1.2. Exponential intermediate region
In the exponential case, one may construct a symmetrizer on $(-\infty,\delta]$ as in the smooth case, then join to essentially arbitrarily weighted frozen-coefficient symmetrizers constructed by WKB, by the same interpolation argument of the previous subsection, weighting the decaying mode sufficiently strongly, as in the proof of Proposition 3.9. The symmetrizer on [x~,+)[$\tilde${x}-$\infty$$\infty$) may then be chosen by a straighforward weighting as in Section 3.4.1.
6.1.3. Turning point at infinity
The case of a turning point near -$\infty$, with shock location taken without loss of generality at $x=0$, reduces essentially to the bounded domain computation of the previous subcases. For, as noted in the treatment of smooth profiles, the same constant symmetrizer used for finite turning points applies also near infinity, up to the value x=Mx=-M, $M\gg 1$ such that $a(-M)=a_{0}$, $a_{0}>0$ small and fixed. Using the same matching arguments as previously, we reduce again to the bounded domain problem of Section 4.2 on [M,0][-M,0] with artifical left boundary condition $u^{+}(0)=u^{-}(0)$, and the argument goes as before.
6.1.4. Coinciding subshock and turning point
In the case that the turning point coincide (or nearly coincide), one must modify the choice of symmetrizer to emphasize the weight of the decaying mode, in order to achieve boundary dissipativity, (3.4). The same issue arises in the constant-coefficient case treated by Kreiss, and can be handled following the approach of [Lemmas 4.1-4.4]K.
Namely, substituting for (6.9) the modified symmetrizer
(6.13) $S^{d}_{\alpha}=\begin{pmatrix}0&1+i\varepsilon\tilde{\gamma}/2\alpha\\ 1-i\varepsilon\tilde{\gamma}/2\alpha&d\end{pmatrix}$.
for $d>0$ sufficiently large and $\alpha>0$ sufficiently small, we find that $S_{\alpha}$ satisfies both interior and boundary dissipativity on a small neighborhood $\tilde{x}-\delta,\tilde{x}$] of the turning point and shock. Note that this is exactly analogous to the choice in [24] of symmetric principal part $D$ and corrector $iF$, $F$ skew, given by
(6.14) $D=\begin{pmatrix}0&1\\ 1&d\end{pmatrix},\qquad F=\begin{pmatrix}0&C\\ -C&0\end{pmatrix}$,
with $d$, $C$ sufficiently large, in the simplest setting of an $s\times s$ Jordan block with $s=2$.
[…]
(6.15) $S_{WKB}:=L^{*}\begin{pmatrix}\theta&0\\ 0&-1\end{pmatrix}L$
generalizing (5.14). For, as noted in [24, Lemma 4.1], there is a unique one-parameter family of symmetrizers $D$ of the principal Jordan block part
$\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$
[…]
6.1.5. Basic discontinuous damping theorem
Collecting the above arguments, we obtain the following basic result, sufficient to treat multi-D stability in (some) physically interesting situations.
Theorem 6.4 (Discontinuous case).
For profiles with a single discontinuity, under the structural assumptions of Theorem 6.2 (in particular, at most a single Airy-type turning point in each separate block), together with linear high-frequency resolvent bound (1.3) for the associated linearized problem, there exists a smooth family of Kreiss symmetrizers for (6.3) with uniform constant $\tilde{\gamma}>0$ for $0<\varepsilon\leq\varepsilon_{0}$ sufficiently small.
Remark 6.5.
As noted in [11, 55], the hypotheses of Theorem 6.4 are satisfied in the oscillatory indermediate region case for all discontinuous monotone decreasing profiles of the Saint Venant equations for inclined shallow-water flow, which occur precisely in the hydrodynamically stable case of Froude number $F<2$. Nonmonotone discontinuous profiles, occurring for $F>2$, fall in the exponential intermediate region case, but do not satisfy the dissipitivity condition and endstates, since this is equivalent to hydrodynamic stability. It would be very interesting to establish convective nonlinear stability as done in the 1-D case in [11], i.e., stability in an appropriately exponentially weighted norm, as observed numerically in [55].
Remark 6.6.
Our treatment of the oscillatory case is not a simple analog of the constant-coefficient Kreiss Theory [24], but a blend of that theory with the rather different set of ideas described in Section 4.2, originating in the 1-D investigations of Rodrigues, Duchêne, Gareaux, Faye, and the author, as described in [50], with motivations also from the spectral investigations of [8, 27, 61, 55].
@@ 7.2. Truncation to finite time
7.3. Absorbing nonlinear sources $f$ and $g$
In the smooth multi-d case, it is straightforward to absorb nonlinear contributions already at the level of symmetrizer estimates, yielding a nonlinear damping bound
(7.7) $\|v(T)\|_{H^{s}_{\alpha}}^{2}\leq Ce^{-\eta T}\|v(0)\|_{H^{s}_{\alpha}}^{2}+C\int_{0}^{T}e^{-\eta(T-t)}\|v(t)\|_{L^{2}_{\alpha}}^{2}\,dt$
corresponding to that in [57, 58], and sufficient to close the estimates as done in these references. In the 1-D case, term $\psi_{t}\bar{w}_{x}$ in $f$ (and, in the discontinous case, multiples of $\psi_{t}$ in $g$) leads to the modified bound
(7.8) $\|v(T)\|_{H^{s}_{\alpha}}^{2}\leq Ce^{-\eta T}\|v(0)\|_{H^{s}_{\alpha}}^{2}+C\int_{0}^{T}e^{-\eta(T-t)}\Big(\|v(t)\|_{L^{2}_{\alpha}}^{2}+|\psi_{t}(t)|^{2}\Big)\,dt$
corresponding to the nonlinear damping estimate obtained in [34, 35, 54] by Kawashima-type energy methods, and is again sufficient to close the nonlinear argument as done in the smooth case in [34, 35] and in the discontinuous case in [54].
In the discontinuous multi-D case, terms $\nabla_{x_{2},\dots,x_{d},t}\psi\bar{w}_{x}$ in $f$ lead, similarly, to
$\|v(T)\|_{H^{s}_{\alpha}}^{2}\leq Ce^{-\eta T}\|v(0)\|_{H^{s}_{\alpha}}^{2}+C\int_{0}^{T}e^{-\eta(T-t)}\Big(\|v(t)\|_{L^{2}_{\alpha}}^{2}+|\nabla_{x_{2},\dots,x_{d},t}\psi(t)|^{2}\Big)\,dt,$
which should be sufficient to close a nonlinear iteration by arguments like those of [54]. However, this last step is an aspect of the discontinuous multi-D case (only) that has yet to be carried out.
@@ 8.1. Extensions and simplification
Our results yield immediately a number of simplifications and extensions for existing work. For example, In the 1-D large-amplitude relaxation case, they recover and somewhat simplify the damping result of [35], removing contortions associated with translating linear algebraic estimators to Friedrichs symmetrizer form. Likewise, in the viscous shock case [57, 58], global multi-D stability is known only for hyperbolic-parabolic systems of “Kawashima type” (roughly, simultaneously symmetrizable) at the endstates of the shock. Yet, high-frequency bounds (1.3) have been shown in the different context of bounded-time vanishing viscosity already in [16, 17] by Kreiss symmetrizer techniques under the much weaker, apparently sharp assumption [16, 17, (H5)] that the endstates be stable as constant solutions, a condition on the dispersion relation of the associated Fourier symbol. As simultaneous symmetrizability was used in [57, 58] only to establish a nonlinear damping estimate, we thus extend immediately the multi-D long-time stability results of [57, 58] to the larger class of systems considered in [16, 17], putting the two theories (long-time vs. small-viscosity) finally on par.
@@ 8.3. The large-amplitude multi-D case
As regards our main motivating example, of large-amplitude multi-dimensional shocks in the Saint Venant model, the preliminary WKB analysis has been already done in [55, §2.4, p. 32] for the linear problem, yielding the high-frequency linearized spectral stability needed for our analysis, and verifying the assumptions made in Theorem 6.2. We note further, as pointed out in [55], that Airy points do appear for waves of all amplitudes, both smooth and discontinuous type, so that the analysis of Theorem 6.2 is indeed needed to obtain nonlinear damping. Combining these ingredients with the linearized resolvent bounds obtained in [25] for arbitrary-amplitude smooth multi-D profiles of general relaxation systems, one obtains the result that spectral stability in the sense of a standard Evans function condition [25, 55] implies linearized and nonlinear stability with standard decay rates matching those of the parabolic case [57, 58, 59]. Evans stability has been demonstrated for all such waves by nonrigorous numerics in [55]. It would be very interesting to convert this to rigorous computer-aided proof.
@@ 9.1. Open problems
A further issue is the complete treatment of nonlinear stability of multi-D discontinuous “subshock”-containing solutions, generalizing 1-D results of [54, 11]. Here, we have presented without giving full details the main ingredient for nonlinear damping of construction of symmetrizers for critical frequencies $\lambda=i/\varepsilon+\gamma$. As mentioned at the end of Section 7.3, another important step, not appearing in the smooth multi-D theory, is the integration in a full nonlinear stability argument of the $\nabla_{x,t}\psi$ terms appearing in the nonlinear source terms $f$ and $g$. Moreover, the requisite low-frequency linear bounds, though accessible to the same inverse Laplace-transform techniques used in the smooth case, have up to now not been carried out: in particular the estimation of $\nabla_{x,t}\psi$ terms. We expect but have not verified that these can be treated in a similar way as in the 1-D analysis of [54]. We regard this, along with systemization/generalization of the rather special arguments presented here, as the next main open problem in the theory.
In short, beyond the various applications described in Section 8, the main contribution of the present work in our view is to take the theory to the threshold of a Kreiss symmetrizer-based treatment of nonlinear damping and full nonlinear stability of large-amplitude discontinuous profiles in multi-D, under the assumptions of noncharacteristic hyperbolicity (A1) and constant multiplity (A2). Where constant multiplicity fails, e.g., some cases in MHD and anisotropic wave motion, further analysis is needed, even in the constant-coefficient case; see related discussion in [43, 18].
Likewise, failure of noncharacteristicity is an important direction for further study, with applications to periodic roll wave solutions [20, 55]. For, as shown in [50], failure of condition (A1) does not imply failure of damping, nor linear equivalence, but only of the general arguments developed above. Given the important example of Saint Venant roll waves [20, 50], it would be very interesting to investigate this case further using other techniques.
A further interesting problem, as described in Remark 4.4 is the treatment of three or more subshocks or systems of size $n>3$ in the 1-D setting, and associated linear algebraic issues. We note that the classical symmetrizers constructed for roll waves in [50] are “sharp”, in the sense that they exist whenever there holds the high-frequency resolvent bound (1.3), for systems of size $n\leq 5$, but apparently not in general. It would be very interesting, therefore, to reconsider this problem from the symmetrizer point of view, a related linear algebraic question, with the hope of obtaining a sharp result also for systems of size $n>5$.